# Compositional Zero-shot Learning via Progressive Language-based Observations

### v1

@@ C Implementation Details
Adapter [6] 49.6 52.7 39.0 22.2 67.8 75.6 53.1 42.0 43.9 38.2 32.2 14.5
LoRA [7] 50.2 52.9 38.6 22.3 69.1 73.8 55.6 43.2 42.8 38.3 31.8 14.2
@@
Table 6: Ablation study on different PEFT strategies in closed-world settings.
D Extra Quantitative Results
D.1 Ablation Study on PEFT
Setting. To optimize the performance of PLO-VLM for CZSL tasks, we focused on enhancing CLIP’s image encoder using two distinct Parameter Efficient Fine-Tuning (PEFT) strategies: Adapter [6] and LoRA [7], the results are shown in Table 6. Below, we detail their integration into the image encoder of the CLIP.
1) Adapter: Adapters are particularly effective for our purpose as they introduce minimal parameters to the model, allowing us to fine-tune CLIP’s image encoder specifically for the CZSL tasks while avoiding overfitting. The Adapter layers are inserted between the transformer blocks of the image encoder, and they work as follows:
$\text{Ada}(\mathbf{h})=\mathbf{h}+\mathbf{U}(\text{ReLU}(\mathbf{D}(\mathbf{h}))),$ (16)
where $\mathbf{D}(\cdot)$ and $\mathbf{U}(\cdot)$ are the downsampling and upsampling projection functions applied to the hidden states $\mathbf{h}$, with ReLU activation promoting non-linearity. The Adapter layers are optimized during training to adapt the pre-trained representations for better handling of the compositionality in CZSL.
2) LoRA: LoRA is applied to the self-attention modules within the CLIP’s image encoder to achieve a fine-tuning of attention mechanisms tailored to novel CZSL tasks. For each attention module, we modify its weight matrix $\bm{W}$ using low-rank updates as follows:
$\text{LoRA}(\mathbf{h})=\mathbf{h}\times(\bm{W}+\bm{\delta W}),$ (17)
where $\bm{\delta W}=\bm{A}\bm{B}^{T}$ is the low-rank update with learnable matrices $\bm{A}$ and $\bm{B}$. These updates are applied during fine-tuning, enabling the encoder to modify its attention patterns without altering the entire pre-trained weight structure.
@@ D.1 Ablation Study on PEFT
These PEFT strategies are incorporated into the image encoder $En_{v}(\cdot)$ by inserting learnable layers into its transformer blocks. This approach allows for fine-tuning on CZSL tasks while leveraging the rich feature extraction capabilities of CLIP’s pre-trained model.
@@
Results. From Table 6, we can observe that both Adapter and LoRA enhance CLIP for CZSL, with varying degrees of efficacy across different datasets. In the MIT-States dataset, LoRA outperforms Adapter in terms of seen, unseen, and AUC metrics, while Adapter demonstrates slightly better performance in HM metric. On the UT-Zappos dataset, LoRA significantly exceeds Adapter, notably in seen and HM metrics. For the C-GQA dataset, the performance of both strategies is closely matched, with Adapter slightly leading in the seen and HM metrics.
D.2 Observation Order in PLO-LLM
@@
E Extra Qualitative Results
@@
Success and Failure Cases. Figure 8 displays the top-$K$ predictions from our PLO-VLM on the MIT-States, UT-Zappos, and C-GQA datasets. Successful cases, highlighted in green, demonstrate precise model predictions, exemplified by instances such as “sliced chicken”. Notably, some failures, marked in red, such as “gray elephant”, do not align with the ground truth (GT), yet still represent existing compositions within the image. These results underscore our model’s capability to identify state-object compositions, affirming its effectiveness in interpreting complex visual scenes despite occasional misalignments with the GT.
Visualization results of PLO-LLM. In Figure 9, we presented visualizations of the top-1 predictions obtained by PLO-LLM, in cases where observations (obs) were utilized or not. All results were from the MIT-States dataset under the closed-world setting. Concurrently, we displayed the sequence of step-by-step observation prompts on the right side of each image. With our meticulously designed prompt, the LLM effectively generated a progression of observation prompts that transitioned from easy to hard. Through the aid of these well-organized observation prompts, we empowered the model with observing capabilities, leading to better holistic understanding and considerable performance gains.
F Potential Negative Societal Impacts.
While the PLO-VLM and PLO-LLM models represent advancements in visual and linguistic AI applications, they may also have unintended societal impacts. The dependence of PLO on language models and vision-language pre-training raises concerns about perpetuating incorrect or biased interpretations of compositions. This could occur if the underlying data or language prompts that inform these models are biased or flawed, leading to misrepresentations and potentially reinforcing stereotypes or inaccuracies.
G Limitations

### v2

@@ Appendix B Observation Cue Generation
Adapter (Houlsby et al., 2019) 49.7 52.8 39.4 22.3 67.8 75.6 53.1 42.0 43.9 38.2 32.2 14.5
LoRA (Hu et al., 2022b) 50.2 52.9 38.6 22.3 69.1 73.8 55.6 43.2 42.8 38.3 31.8 14.2
MoA (Zheng et al., 2024) 51.6 53.7 40.2 23.4 70.3 75.8 55.3 43.6 44.7 38.1 33.0 14.9
@@
Table 5. Ablation study on different PEFT strategies in closed-world settings.
In addition, this process adheres to the Chain-of-Thought concept (Kojima et al., 2022; Zhang et al., 2022b; Li et al., 2025), generating observations step by step from easy to hard through the utilization of a special prompt “Let’s observe it step by step!”.
Filter strategies. When the query contains an unreasonable composition category, it is always difficult for LLM to generate the corresponding (or the same) format as those in-context examples given by the instruction, e.g., LLM may reply “For this concept, which is nonsensical and not realistically possible”. We have used two filter strategies for those unreasonable composition queries:
  • Unreasonable observation replace. We apply strict format matching (item numbers and the prefix “a photo of”) to filter out unreasonable categories and replace them with hard prompts. Specifically, we use the “a photo of [state] [object]” to replace the generated observations/descriptions of the unreasonable composition category.
  • Unfeasible categories filter. During inference, existing state-of-the-art methods also define the list of unfeasible categories and filter out unfeasible categories during inference. In our experiments, we followed the same experimental setup as prior work (Mancini et al., 2021).
Note that our strategy DOES NOT pose the risk of information leak. The reasons are: 1) We use exactly the same setting following previous works (Menon and Vondrick, 2023; Li et al., 2023b). Concretely, the state and object categories are pre-defined and remain unchanged. At testing, the same state and object categories from training are used, but with novel compositions appearing. Since CZSL is a classification task, complete composition categories must be provided at test time. Hence, for unseen compositions in testing, we deploy LLMs to generate observation sequences (hard prompts) in the same way. 2) For hard prompt generation, we only provide the state and object category information, and DO NOT offer any image information.
Appendix C Extra Quantitative Results
C.1. Cross-Domain Evaluation
Setting. DFSP(i2t) (Lu et al., 2023a), PLO-VLM, and PLO-LLM on the MIT-States (Isola et al., 2015) dataset and tested their cross-domain performance on the C-GQA dataset (Yu and Grauman, 2014). We selected categories corresponding to the states and objects in MIT-States to ensure a consistent assessment. In addition, we also conducted cross-domain tests on the MIT-States dataset in reverse.
Results. As shown in Table 3, PLO-LLM exhibits superior performance on both C-GQA and MIT-States datasets, particularly in the accuracy of unseen categories (32.1% vs. 28.7% and 43.6% vs. 42.4%) and consistently outperforms in HM and AUC scores. This enhancement stems from the GPT-generated sequence of observation cues in PLO-LLM, which operates independently of the visual modality, thereby bolstering cross-domain robustness.
C.2. Comparison with the LLM-based Models
To prove the effectiveness of step-by-step observations, following (Menon and Vondrick, 2023), we added cues generated by LLM in the class label (CLS-DES). The results are shown in Table 4. It can be seen that our step-by-step fused strategy can get higher performance than cue comparison at the class-level. This phenomenon might be attributed to the necessity of one-step recognition for different cues at the class level. Unlike the one-step identification of composition, progressive observation allows for a greater focus on salient features, thereby enabling more precise predictions.
C.3. Ablation Study on PEFT
@@
Setting. To optimize the performance of PLO-VLM for CZSL tasks, we focused on enhancing CLIP’s image encoder using two distinct Parameter Efficient Fine-Tuning (PEFT) strategies: Adapter (Houlsby et al., 2019) and LoRA (Hu et al., 2022b). Besides, we also conducted MoA strategy (Zheng et al., 2024) on our PLO-VLM, which reinforces both CLIP’s image encoder and text encoder in the same manner as (Zheng et al., 2024). These PEFT strategies are incorporated into the image encoder $En_{v}(\cdot)$ by inserting learnable layers into its transformer blocks. This approach allows for fine-tuning on CZSL tasks while leveraging the rich feature extraction capabilities of CLIP’s pre-trained model. The results are shown in Table 5.
@@
Results. From Table 5, we can observe that all three PEFT strategies can enhance CLIP for CZSL, with varying degrees of efficacy across different datasets. Benefiting from the more parameters that can be fine-tuned, the MoA strategy of enhancing both the image encoder and text encoder can make PLO-VLM achieve the best performance over all datasets. As for mere enhancement in image encoder, LoRA outperforms Adapter in terms of both seen and unseen metrics, while Adapter demonstrates slightly better performance in HM metric in the MIT-States dataset. On the UT-Zappos dataset, LoRA significantly exceeds Adapter, notably in seen and HM metrics. For the C-GQA dataset, the performance of both strategies is closely matched, with Adapter slightly leading in the seen and HM metrics.
Figure 9. Distribution of samples where state or object is first observed in PLO-VLM on the test set of MIT-States.
 Backbone Method MIT-States
S U HM AUC
ViT-B/32 DFSP (Lu et al., 2023a) 36.7 43.4 29.4 13.2
PLO-VLM 41.1 44.2 31.3 14.8
PLO-LLM 44.2 47.9 34.3 17.4
ViT-B/16 DFSP (Lu et al., 2023a) 39.6 46.5 31.5 15.1
PLO-VLM 43.9 46.6 34.1 17.0
PLO-LLM 41.2 44.7 31.4 14.8
ViT-L/14 DFSP (Lu et al., 2023a) 46.9 52.0 37.3 20.6
PLO-VLM 49.7 52.8 39.4 22.3
PLO-LLM 49.6 53.2 39.0 21.9
Table 6. Ablation study on different CLIP backbones.
C.4. Ablation Study on Network Architectures
We further examined the influence of replacing CLIP backbones in our PLO-VLM and PLO-LLM. All results are reported in Table 6. Those consistent performance gains compared to the previous SOTA method reaffirmed the efficacy and robustness of the proposed methodology. For all other experiments, we chose ViT-L/14 by default.
C.5. Observation Order in PLO-VLM
@@
 $K$-time MIT-States UT-Zappos C-GQA
S U HM AUC S U HM AUC S U HM AUC
1 49.7 52.8 39.4 22.3 67.8 75.6 53.1 42.0 43.9 38.2 32.2 14.5
2 49.6 53.0 39.4 22.3 68.1 75.3 53.4 42.1 43.7 38.1 32.0 14.4
3 49.8 52.7 39.3 22.2 68.2 75.5 53.0 42.2 43.8 38.1 32.1 14.4
4 49.7 52.7 39.3 22.2 68.0 75.5 53.2 42.1 43.5 38.2 32.0 14.3
5 49.9 52.8 39.4 22.3 67.8 75.8 52.9 42.0 43.7 38.3 32.1 14.4
AvgAvg 49.74 52.80 39.36 22.26 67.98 75.54 53.12 42.08 43.72 38.18 32.08 14.40
StdevStdev 0.11 0.12 0.05 0.05 0.18 0.18 0.19 0.08 0.15 0.08 0.08 0.07
Table 7. The $K$-times experiment of PLO-VLM in closed-world setting.
C.6. Observation Order in PLO-LLM
  Order MIT-States
S U HM AUC
(1, 2, 3) 49.6 53.2 39.0 21.9
(1, 3, 2) 48.8 52.9 38.7 21.8
(2, 1, 3) 48.6 53.1 38.3 21.6
(2, 3, 1) 48.8 53.3 38.4 21.7
(3, 1, 2) 49.0 53.1 38.2 21.6
(3, 2, 1) 48.8 52.3 38.5 21.4
Table 8. Ablation study on cue orders in PLO-LLM.
We also conducted ablation study on the order of cues in Table 8, revealing that the sequence slightly impacts performance. The last cue is fixed as “a photo of state object” and we change the order of the previous three cues generated by LLM. As seen, when cues follow an easy-to-difficult progression, PLO-LLM achieves better performance. Conversely, a reverse order, from difficult to easy, results in the lowest performance. This suggests that starting with simpler cues allows the model to gradually build a foundational understanding.
C.7. Robustness of PLO
To analyze the robustness of our PLO, we provided the performance deviations of PLO-VLM by $K$-times experiments with the same parameters. The results are shown in Table 7. As seen, PLO-VLM can stably outperform existing methods (cf., Table 1).
C.8. Ablation Study on the Hard Prompt and Soft Prompt
 $\lambda$ MIT-States
S U HM AUC
0.0 40.3 42.9 30.1 14.0
0.3 47.1 52.9 35.4 20.7
0.5 48.3 52.9 37.4 21.4
0.7 49.6 53.2 39.0 21.9
1.0 49.0 52.9 38.5 21.6
Table 9. Ablation study on different fusion weights of the hard and soft prompts.
Hard prompts generated by LLM can provide structured guidance for refining visual features, but also face limitations such as potential misalignment with CLIP’s semantic space and datasets’ underlying distribution. Soft prompts offer flexibility, enabling context-aware adjustments and leveraging CLIP’s broad knowledge for task-specific applications with learnable tokens (Zhou et al., 2022b). As seen in Table 9, we studied the effect of fusion weight ($\lambda$ for hard prompts and $1-\lambda$ for soft prompts). From the results we can observe that combining hard and soft prompts together leads to the best performance. $\lambda$ is set to be 0.7 by default.
Figure 10. Top-$K$ predictions on for randomly selected cases from three datasets. The top and bottom rows show the results of closed-world settings, respectively. Correct and incorrect predictions are highlighted in green and red, respectively.
Figure 11. Top-$K$ predictions on for randomly selected cases from MIT-States. The top and bottom rows show the results of open-world settings, respectively. Correct and incorrect predictions are highlighted in green and red, respectively.
Appendix D Extra Qualitative Results
D.1. Success and Failure Cases
@@
Figure 10 displays the top-$K$ predictions from our PLO-VLM on the MIT-States, UT-Zappos, and C-GQA datasets under closed-world setting. Successful cases, highlighted in green, demonstrate precise model predictions, exemplified by instances such as “sliced chicken”. Notably, some failures, marked in red, such as “gray elephant”, do not align with the ground truth (GT), yet still represent existing compositions within the image. These results underscore our model’s capability to identify state-object compositions, affirming its effectiveness in interpreting complex visual scenes despite occasional misalignments with the GT. We also provided top-$K$ predictions of PLO-VLM in the open-world settings in Figure 11. Our progressive observation strategy enables a thorough grasp of compositions, which is particularly effective in recognizing unseen compositions by bridging the gap between base and novel categories. Notably, even when our model’s top-1 prediction is not exactly matched, it still accurately predicts the state primitive, and the object primitive also appears to be reasonable (e.g., muddy stream). These results underscore PLO’s effectiveness in capturing pertinent visual cues and enabling comprehensive composition understanding.
D.2. Visualization Analysis of PLO-VLM
Figure 12. Confusion matrix results for different models on test samples labeled as “mashed potato”.
As mentioned above, we emphasize that identifying a “mashed object” instead of directly determining the entire composition leads to better performance. Yet, not all state-first methods ensure precise object classification. Variations in mashing degrees for potatoes and bananas present challenges. Lightly mashed samples are distinguishable by potato chunks versus banana stickiness. However, heavily mashed samples may confound even expert annotators.
To validate the effectiveness of PLO-VLM, we visualized confusion matrices of different models for testing samples labeled as “mashed potato” in Figure 12. We outline the following key points: 1) As for the “State First” and “Dynamic Observation” cases, distinguishing between similar states such as “whipped” and “crushed” poses a significant challenge. Such similar states represent more intense (whipped) or milder (crushed) variations of the “mashed” state, leading to potential errors in state prioritization. Note that even humans find it difficult to distinguish these similar states. 2) Despite the aforementioned challenges, PLO-VLM outperforms DFSP by a clear margin, indicating PLO-VLM’s capability to recognize such nuanced differences. 3) The “Object First” strategy might mistakenly identify objects, e.g., “garlic”, due to garlic often being presented in a mashed form and sharing yellow and white colors in different states. PLO-VLM addresses this by first categorizing “mashed object” and then analyzing state-specific traits like texture and viscosity.
Appendix E Discussion of two PLO Variants
Although the observation steps of PLO-VLM and PLO-LLM are different, they have a common motivation and common progressive observation process.
  • Common Motivation: Both methods are designed to address primitive-conditioned variances. As detailed in Sec. 1, our motivation starts with the concern that PLO-VLM’s two-step observation method based on visual features may overfit specific datasets. Subsequently, we further introduce PLO-LLM to enhance robustness through a flexible, multi-step observation process.
  • Common Progressive Observation Process: Both models employ a sequential refinement of visual features, albeit through different strategies (two-step PLO-VLM vs.multi-step in PLO-LLM). This process utilizes the cross-attention mechanism to leverage cues from preceding steps for progressive observation, as demonstrated in Eq. (4) and Eq. (7).
Appendix F Limitations